OpenStudy (anonymous):

y = x + 4 –3x + 2y = 8 Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. I really do need the help

OpenStudy (anonymous):

\[y-mx+b\]Where \(m\) is the slope and \(b\) is the value of y at the y-intercept.

OpenStudy (anonymous):

I mean \(y=mx+b\)

OpenStudy (wwe123):

x=0, y = 4

OpenStudy (wwe123):

y = x + 4 –3x + 2(x + 4) = 8 -3x+2x+8=8 -x=0 x=0 y= x + 4 y=(0) + 4 y=4